# Towards Robust Infrared Small Target Detection: A Feature-Enhanced and Sensitivity-Tunable Framework

### v2

@@ 3.3 Edge enhancement difficulty mining loss
In the task of SIRST detection, the small target area and lack of intrinsic features make it difficult to accurately locate in the image. Existing studies have shown that paying attention to edge information in this task can bring significant performance improvements [31, 33]. Therefore, on the basis of the characteristics of infrared small targets, we design an EEDM loss. This loss can effectively guides the network to pay more attention to edge features and challenging target regions during training, thereby enhancing the model’s boundary awareness and improving its ability to learn from challenging samples.
The EEDM loss consists of two components: edge pixel enhancement and difficult pixel mining. Edge pixel enhancement is performed preferentially. Specifically, the edge contour of the infrared small target is first extracted from the truth label obtain the edge map $e$, which is then weighted using corresponding edge weight coefficient $w$. Secondly, the binary cross-entropy loss [38] is employed to compute the loss value for each pixel, thereby constructing the matrix ${L_{BCE}}$. Finally, the edge weighting matrix is used to apply a weighted calculation to the loss matrix ${L_{BCE}}$, resulting in the weighted loss value ${L^{*}}$. The formulas are as follows:
${L_{BCE}}={y_{i}}\log({\hat{y}_{i}})+(1-{y_{i}})\log(1-{\hat{y}_{i}})$ (1)
@@
where ${y_{i}}$ denotes the true label of the $i_{th}$ pixel, ${\hat{y}_{i}}$ denotes the probability of the $i_{th}$ pixel being predicted as the target, and $e^{\prime}$ denotes the edge weighting matrix.
The difficult pixel mining is performed subsequently. Specifically, first, the loss values corresponding to each pixel in ${L^{*}}$ are sorted. Secondly, pixels with loss values greater than or equal to the difficult pixel mining ratio are selected as the set of difficult pixels ${L_{{\rm{hard}}}}$. Finally, the average loss of the difficult pixels is computed to obtain the final loss value. The formulas are as follows:
@@
where $p$ denotes the difficult pixel mining ratiol and ${L_{EEDM}}$ denotes the final output.
@@ 3.4 Adjustable sensitivity strategy
In the infrared small target detection task, the network outputs a probability mask, in which the value of each pixel denotes its confidence of belonging to the target. Existing studies typically use a single threshold (0.5) to separate targets from the background [4, 32]. However, we believe that the inherent idea of directly using a single threshold cannot fully mine and utilize the information in the probability mask. Therefore, we propose an AS strategy aimed at deeply mining the information in the probability mask. Based on this, we innovatively define the concepts of strong targets and weak targets. Strong targets refer to those that are easy for the network to detect and have high confidence. These targets typically display more distinct features in the image, enabling the network to quickly identify them through learning. For strong targets, our goal is to achieve complete and precise segmentation. In contrast, weak targets refer to those that are difficult for the network to detect and have lower confidence. These targets may be challenging for the network to recognize due to their shape or the complexity of the background. For weak targets, our goal is to identify them to meet the needs of practical scenarios.
Figure 3: Some samples from multiple datasets.
@@
Through careful observation of the probability mask generated by deep learning methods in this task, we interestingly found that, influenced by data ambiguity and the network’s feature extraction mechanism, the network’s predicted probability mask for the target region exhibits a Gaussian-like distribution. Specifically, the probability is higher in the center region and gradually decreases towards the edge, and finally approaches zero. Based on this observation and to better detect strong and weak targets, we design two thresholds th1 and th2 ($th1>th2$) to adjust the model’s responsiveness to regions with different confidence levels. Specifically, when the network’s confidence in a region is higher than th1, the region is defined as a strong target region. In this case, the network demonstrates high detection capability and can provide relatively accurate segmentation results. Additionally, when the confidence is between th1 and th2, the region is defined as a weak target region. For these weak target regions, although the network’s detection accuracy is low, reasonable post-processing can still help identify partial features of these targets to meet the detection needs of weak target detection in practical scenarios.
@@
The specific process of the AS strategy is shown in fig. 2, first, we separate the segmentation result ${M_{th2}}$ of th2 into multiple binary images $M_{th2}^{s},\;s=1,2,\ldots n$ with a single target region according to a single connected region. Secondly, these binary images $M_{th2}^{s},\;s=1,2,\ldots n$ are compared with the segmentation results ${M_{th1}}$ of th1 in turn to determine whether there are overlapping regions. For $M_{th2}^{s}$ with overlapping regions, we lose it. For the same target region, the fineness in ${M_{th1}}$ is better than that in ${M_{th2}}$. Thirdly, we extract the centroid of $M_{th2}^{s}$ where there is no overlapping region. Finally, the extracted centroid points are injected into ${M_{th1}}$ to generate the final segmentation result map. The highlighted weak target region is added to ${M_{th1}}$ in the form of centroid points without affecting the fine segmentation of the strong target region in ${M_{th1}}$. It is worth noting that this strategy exhibits good generality, as it can be seamlessly integrated into existing networks as a post-processing step without requiring any modifications to the original structure.
@@ 4.1 Dataset
Four SIRST datasets (W-IRSTD dataset [39], NUDT-SIRST [29], NUAA-SIRST [25], and IRSTD-1k [33]) are used for experiments. These datasets are introduced as follows:
1) W-IRSTD dataset: The dataset contains 9,000 images, which mainly come from seven public datasets: SIRST-V2 [40], IRSTD-1k [33], IRDST [41], NUDT-SIRST [29], NUDT-SIRST-sea [42], NUDT-MIRSDT [43] and Anti-UAV410 [5]. fig. 3(d) shows some samples from the W-IRSTD dataset. The dataset covers multiple observation perspectives, including land-based, air-based, and space-based. In addition, it includes diverse target types such as extended targets, spot targets, and point targets, spans multiple spectral bands including short-wave infrared, long-wave infrared, and near-infrared, and provides various resolutions such as 256×256, 512×512, and 1024×1024. In the experiment, 7200 samples are used as the training set and 1800 samples are used as the test set.
2) NUAA-SIRST dataset: The real dataset consists of 427 infrared images captured from various scenes, with approximately 10% of the images containing multiple targets. In the experiment, we divide the dataset into training set and test set in a ratio of 213:214. fig. 3(a) shows some samples from the NUAA-SIRST dataset.
3) NUDT-SIRST dataset: The synthetic dataset covers five background types: city, field, highlight, ocean and cloud. About 37% of the images contain two or more targets, and about 32% of the targets are distributed outside the top 10% of the image brightness values. In the experiment, we divide the dataset into training set and test set in a ratio of 663:664. fig. 3(b) shows some samples from the NUDT-SIRST dataset.
4) IRSTD-1k dataset: The real dataset consists of 1001 infrared images with a resolution of 512×512 pixels. It contains small targets of various types, including drones, creatures, ships, and vehicles, with diverse spatial distributions. Meanwhile, the dataset covers a wide range of scenes, such as sky, ocean, and land. In the experiments, we divide the dataset into training set and test set in a ratio of 801:200. fig. 3(c) shows some samples from the IRSTD-1k dataset.
@@ 4.2 Experimental settings
1) Experimental environment and parameter settings. The experimental environment is an Ubuntu 18.04 operating system, and the GPU is an RTX 3090 24 GB. The epoch is set to 300 and the learning rate is set to $5e^{-4}$.
@@
Scheme $IoU$ $P_{d}$ $F_{a}$$10^{-6}$) ↓ Score
SoftIoU Loss [44] 44.08 53.69 12.39 48.89
Dice Loss [45] 50.15 60.09 9.09 55.12
BCE Loss [38] 55.44 59.35 7.02 57.40
@@
where $N$ denotes the number of images in the test set. $TP$ denotes the small target area that the model correctly predicts. $T$ denotes the small target area in the true label. $P$ denotes the small target area predicted by the model. ${T_{correct}}$ denotes the number of correctly predicted targets. ${T_{All}}$ denotes the true target number. ${P_{false}}$ denotes the number of incorrectly predicted target pixels. ${P_{all}}$ denotes the total number of pixels in the image. $\alpha$ is set to 0.5. In addition, $F_{a}$ has a constraint effect. When $F_{a}$ is less than $1e^{-4}$, the final result is considered valid.
@@ 4.3 Effect verification of the EEDM loss
To further verify the effect of the proposed EEDM loss, we conduct a detailed experimental exploration, including comparative with other common losses [38, 44, 45] and evaluation of hyperparameter stability. The experiments are conducted on the W-IRSTD dataset. MSDA-Net is selected as the basic model.
1) Comparison with other loss functions: To evaluate the effectiveness of the proposed EEDM Loss, we conduct comparative experiments against commonly used loss functions, including binary cross-entropy Loss (BCE Loss) [38], soft intersection over union Loss (SoftIoU Loss) [44] and Dice Loss [45]. To comprehensively evaluate the effectiveness of the proposed EEDM Loss, we further introduce two ablation variants for comparison: the edge enhancement loss (EE Loss), which contains only the edge enhancement part, and the difficult mining loss (DM Loss), which contains only the difficult pixel mining part. This design enables a quantitative analysis of the respective contributions of edge enhancement and difficult sample mining to the overall detection performance.
From table 2, compared with other losses, the proposed EEDM loss results in more obvious improvements in both $IoU$ and $P_{d}$, verifying its superiority in infrared small target detection tasks. In contrast, the widely used SoftIoU Loss exhibits relatively inferior performance on the complex and diverse W-IRSTD dataset. At the same time, compared with BCE Loss, EE Loss achieves improvements of 1.48% in $IoU$ (from 55.44 to 56.26), 1.87% in $P_{d}$ (from 59.35 to 60.46), and 1.67% in Score (from 57.40 to 58.36), indicating that incorporating edge features during training can effectively enhance detection performance. In addition, compared with BCE Loss, DM Loss achieves improvements of 1.23% in $IoU$ (from 55.44 to 56.12), 5.2% in $P_{d}$ (from 59.35 to 62.45), and 3.19% in Score (from 57.40 to 59.29), indicating the effectiveness of difficult pixel mining in this task. From fig. 4, the ROC curve of the model trained with EEDM Loss consistently stays above the others, achieving the highest AUC, which indicates superior performance across all thresholds compared to other loss functions.
Figure 5: Performance analysis of the hyperparameters $w$ and $p$ on the W-IRSTD dataset. From left to right, the figures show the model performance in terms of $IoU$, $P_{d}$, $F_{a}$, and Score under different combinations of $(w,p)$. Red indicates the optimal value in each metric.
@@
2) Validation of hyperparameter stability: To comprehensively evaluate the performance of the proposed EEDM Loss, we conduct detail experiments on two key hyperparameters (the edge weight coefficient $w$ and the difficult pixel mining ratio $p$) to investigate the model’s sensitivity and robustness to different hyperparameter settings. Specifically, the edge weight coefficient $w$ is set to 1, 3, 4, 5, 7, and the difficult pixel mining ratio $p$ is set to 0.1, 0.3, 0.5, 0.7, 0.9.
From fig. 5, the proposed EEDM Loss can maintain relatively stable performance across different hyperparameter configurations. This indicates that the loss function is robust to the hyperparameters $w$ and $p$, as it consistently enhances model performance within a reasonable range of these hyperparameters. Based on the experimental results, the metrics Score reaches the optimal value when $w$ is 4 and $p$ is 0.5. Specific analysis shows that when the $w$ is too small, the model pays insufficient attention to edge information, leading to decreases in $IoU$ and $P_{d}$. When $w$ is too large, the model will overly rely on edge features, thereby neglecting semantic and other critical features, which negatively impacts the overall detection performance. It is worth noting that when $w$ is extremely reduced to 1, the EEDM Loss degenerates into the DM Loss. Meanwhile, when the $p$ is too small, the network tends to excessively focus on difficult pixels during training, which may hinder its generalization ability and lead to degraded performance. Conversely, when $p$ is too large, the mining of difficult pixels will be insufficient, thereby weakening the model’s adaptability in complex scenarios. It is worth noting that when $p$ increases to 1, the EEDM Loss degenerates into the EE Loss.

### v3

@@ 3.3 Edge enhancement difficulty mining loss
In SIRST detection, the small target area and lack of intrinsic features make it difficult to accurately locate in the image. Existing studies have demonstrated that paying attention to edge information in this task can bring significant performance improvements [zhao2023gradient, zhang2022isnet]. Therefore, on the basis of the characteristics of infrared small targets, we design an EEDM loss. This loss can effectively guide the network to pay more attention to edge features and challenging target regions during training, thereby enhancing the model’s boundary awareness and improving its ability to learn from challenging samples. The workflow of the EEDM loss is shown in Fig. 2.
The EEDM loss consists of two components: edge pixel enhancement and difficult pixel mining. Edge pixel enhancement is performed preferentially. Specifically, the pixel-wise loss values are first computed using the binary cross-entropy loss [li2024rediscovering], yielding the loss matrix $L_{BCE}$}. At the same time, we perform edge extraction on the true label to obtain the target edge map. Based on this, an edge weighting coefficient $w$ is introduced to weight the target edge map and the loss matrix $L_{BCE}$}, resulting in the weighted loss matrix $L^{*}$}. The formulas are as follows:
${L_{BCE}}=-[{y_{i}}\log({\hat{y}_{i}})+(1-{y_{i}})\log(1-{\hat{y}_{i}})]$ (1)
@@
where $y_{i}$} denotes the true label of the $i_{th}$ pixel, $\hat{y}_{i}$} denotes the probability of the $i_{th}$ pixel being predicted as the target, $e$ denotes the target edge map, $w$ denotes the edge weighting coefficient, and $e^{\prime}$ denotes the edge weighting map.
The difficult pixel mining is performed subsequently. Specifically, first, the loss values corresponding to each pixel in $L^{*}$} are sorted. Secondly, pixels with loss values greater than or equal to the difficult pixel mining ratio are selected as the set of difficult loss values $L_{{\rm{hard}}}$}. Finally, the average of the difficult loss values is computed to obtain the final loss value. The formulas are as follows:
@@
where $\text{sort}(\cdot)$ denotes the sort operation, $L_{sorted}$} denotes the sorted sequence of all pixel-wise loss values, $N$ denotes the total number of pixel-wise loss values, $p$ denotes the difficult pixel mining ratio and $L_{EEDM}$} denotes the final output.
@@
Figure 2: Illustration of Edge enhancement difficulty mining (EEDM) loss.
Algorithm 2 FEST Framework (Inference phase)
1:Resize set $\{r_{1},r_{2},r_{3}\}$; Batch infrared images $I_{i=1}^{{N_{b}}}$; Networks $\varphi=\{{\varphi_{r1}},{\varphi_{r2}},{\varphi_{r3}}\}$
2:Predicted mask $\{M\}_{i=1}^{{N_{b}}}$
3:Initialize: Give the trained weights to $\varphi_{r1}$, $\varphi_{r2}$, $\varphi_{r3}$.
4:Step 1: For each batch of input images $\{I\}_{i=1}^{{N_{b}}}$, resize inputs into $\{{I_{r1}},{I_{r2}},{I_{r3}}\}_{i=1}^{{N_{b}}}$
5:Step 2: Feed each resized infrared image batch $\{\{{I_{r1}}\}_{i=1}^{{N_{b}}},\{{I_{r2}}\}_{i=1}^{{N_{b}}},\{{I_{r3}}\}_{i=1}^{{N_{b}}}\}$ into the corresponding networks to obtain prediction maps $\{\{{P_{r1}}\}_{i=1}^{{N_{b}}},\{{P_{r2}}\}_{i=1}^{{N_{b}}},\{{P_{r3}}\}_{i=1}^{{N_{b}}}\}$:
$\{{P_{r1}}\}_{i=1}^{{N_{b}}}\;\leftarrow\;{\varphi_{r1}}(\{{I_{r1}}\}_{i=1}^{{N_{b}}})$
$\{{P_{r2}}\}_{i=1}^{{N_{b}}}\;\leftarrow\;{\varphi_{r2}}(\{{I_{r2}}\}_{i=1}^{{N_{b}}})$
$\{{P_{r3}}\}_{i=1}^{{N_{b}}}\;\leftarrow\;{\varphi_{r3}}(\{{I_{r3}}\}_{i=1}^{{N_{b}}})$
6:Step 3: Apply the multi-model fusion strategy on the prediction maps $\{\{{P_{r1}}\}_{i=1}^{{N_{b}}},\{{P_{r2}}\}_{i=1}^{{N_{b}}},\{{P_{r3}}\}_{i=1}^{{N_{b}}}\}$ to obtain the fused map $P_{fuse}$}:
${P_{fuse}}\;\leftarrow\;Mean(Re(\{{P_{r1}}\}_{i=1}^{{N_{b}}}),\;Re(\{{P_{r2}}\}_{i=1}^{{N_{b}}}),\;Re(\{{P_{r3}}\}_{i=1}^{{N_{b}}}))$
7:Step 4: Based on the prediction maps $\{{P_{r1}}\}_{i=1}^{{N_{b}}}$, $\{{P_{r2}}\}_{i=1}^{{N_{b}}}$, $\{{P_{r3}}\}_{i=1}^{{N_{b}}}$ and the fused map $P_{fuse}$}, apply the proposed AS strategy to generate the final detection mask $\{M\}_{i=1}^{{N_{b}}}$
8:Return: $\{M\}_{i=1}^{{N_{b}}}$
@@ 3.4 Adjustable sensitivity strategy
In the small target segmentation task, the network outputs a probability mask, where each pixel value denotes its confidence of belonging to the target. Existing studies typically use a single threshold (0.5) to separate targets from the background [yu2024easy, zhao2024multi]. However, we believe that the inherent idea of directly using a single threshold cannot fully mine and utilize the information in the probability mask. Therefore, we propose an AS strategy aimed at deeply mining the information in the probability mask. Based on this, we innovatively define the concepts of strong targets and weak targets. Strong targets refer to those that are easy for the network to detect and have high confidence. These targets typically display more distinct features in the image, enabling the network to quickly identify them through learning. For strong targets, our goal is to achieve complete and precise segmentation. In contrast, weak targets refer to those that are difficult for the network to detect and have lower confidence. These targets may be challenging for the network to recognize because of their shape or the complexity of the background. For weak targets, our goal is to identify them to meet the needs of practical scenarios.
@@
Through careful observation of the probability mask generated by deep learning methods in this task, we interestingly find that, influenced by data ambiguity and the network’s feature extraction mechanism, the network’s predicted probability mask for the target region exhibits a Gaussian-like distribution. Specifically, the probability is higher in the center region and gradually decreases toward the edge, and finally approaches zero. Based on this observation and to better detect strong and weak targets, we design two thresholds $th1$ and $th2$ ($th1>th2$) to adjust the model’s responsiveness to regions with different confidence levels. Specifically, when the network’s confidence in a region is higher than $th1$, the region is defined as a strong target region. In this case, the network demonstrates high detection capability and can provide relatively accurate segmentation results. Additionally, when the confidence is between $th1$ and $th2$, the region is defined as a weak target region. For these weak target regions, although the network’s detection accuracy is low, reasonable post-processing can still help identify partial features of these targets to meet the detection needs of weak target detection in practical scenarios.
@@
The specific process of the AS strategy is shown in Fig. 3. First, we separate the segmentation result $M_{th2}$} of $th2$ into multiple binary images $M_{th2}^{s},\;s=1,2,\ldots n$ with a single target region according to a single connected region. Secondly, these binary images $M_{th2}^{s},\;s=1,2,\ldots n$ are compared with the segmentation results $M_{th1}$} of $th1$ in turn to determine whether there are overlapping regions. For $M_{th2}^{s}$ with overlapping regions, we lose it. For the same target region, the fineness in $M_{th1}$} is better than that in $M_{th2}$}. Thirdly, we extract the centroid of $M_{th2}^{s}$ where there is no overlapping region. Finally, the extracted centroid points are injected into $M_{th1}$} to generate the final segmentation result map. The highlighted weak target region is added to $M_{th1}$} in the form of centroid points without affecting the fine segmentation of the strong target region in $M_{th1}$}. It is worth noting that this strategy exhibits good generality, as it can be seamlessly integrated into existing networks as a post-processing step without requiring any modifications to the original structure.
@@ 4.1 Dataset
Four SIRST datasets (W-IRSTD dataset [li2024icpr], NUDT-SIRST [li2022dense], NUAA-SIRST [dai2021asymmetric], and IRSTD-1k [zhang2022isnet]) are used for the experiments. These datasets are introduced as follows:
1) W-IRSTD dataset: The dataset contains 9,000 images, which mainly come from seven public datasets: SIRST-V2 [dai2023one], IRSTD-1k [zhang2022isnet], IRDST [sun2023receptive], NUDT-SIRST [li2022dense], NUDT-SIRST-sea [wu2023mtu], NUDT-MIRSDT [li2023direction] and Anti-UAV410 [huang2023anti]. Fig. 4(d) shows some samples from the W-IRSTD dataset. The dataset covers multiple observation perspectives, including land-based, air-based, and space-based. In addition, it includes diverse target types such as extended targets, spot targets, and point targets, spans multiple spectral bands including short-wave infrared, long-wave infrared, and near-infrared, and provides various resolutions such as 256×256, 512×512, and 1024×1024. In the experiment, 7200 samples are used as the training set and 1800 samples are used as the test set.
2) NUAA-SIRST dataset: The real dataset consists of 427 infrared images captured from various scenes, with approximately 10% of the images containing multiple targets. In the experiment, we divide the dataset into a training set and a test set at a ratio of 213:214. Fig. 4(a) shows some samples from the NUAA-SIRST dataset.
3) NUDT-SIRST dataset: The synthetic dataset covers five background types: city, field, highlight, ocean and cloud. Approximately 37% of the images contain two or more targets, and approximately 32% of the targets are distributed outside the top 10% of the image brightness values. In the experiment, we divide the dataset into a training set and a test set at a ratio of 663:664. Fig. 4(b) shows some samples from the NUDT-SIRST dataset.
4) IRSTD-1k dataset: The real dataset consists of 1001 infrared images with a resolution of 512×512 pixels. It contains various types of small targets, including drones, creatures, ships, and vehicles, with diverse spatial distributions. Meanwhile, the dataset covers a wide range of scenes, such as the sky, ocean, and land. In the experiments, we divide the dataset into a training set and a test set at a ratio of 801:200. Fig. 4(c) shows some samples from the IRSTD-1k dataset.
@@ 4.2 Experimental settings
1) Experimental environment and parameter settings. The experimental environment is an Ubuntu 18.04 operating system, and the GPU is an RTX 3090 24 GB. The number of training epochs is set to 300 and the learning rate is set to $5e^{-4}$.
Figure 4: Some samples from multiple datasets.
@@
Scheme IoUIoU $P_{d}$ $F_{a}$$10^{-6}$) ↓ Score
SoftIoU Loss [rahman2016optimizing] 44.08 53.69 12.39 48.89
Dice Loss [li2019dice] 50.15 60.09 9.09 55.12
BCE Loss [li2024rediscovering] 55.44 59.35 7.02 57.40
@@
where $N$ denotes the number of images in the test set. $TP$ denotes the small target area that the model correctly predicts. $T$ denotes the small target area in the true label. $P$ denotes the small target area predicted by the model. $T_{correct}$} denotes the number of correctly predicted targets. $T_{All}$} denotes the true target number. $P_{false}$} denotes the number of incorrectly predicted target pixels. $P_{all}$} denotes the total number of pixels in the image. $\alpha$ is set to 0.5. In addition, $F_{a}$ has a constraint effect. When $F_{a}$ is less than $10^{-4}$, the final result is considered valid.
@@ 4.3 Effect verification of the EEDM loss
To further verify the effect of the proposed EEDM loss, we conduct extensive experiments, including comparisons with other common losses [li2024rediscovering, rahman2016optimizing, li2019dice] and an evaluation of hyperparameter stability. The experiments are conducted on the W-IRSTD dataset. MSDA-Net is selected as the basic model.
1) Comparison with other loss functions: To evaluate the effectiveness of the proposed EEDM loss, we conduct comparative experiments against commonly used loss functions, including binary cross-entropy (BCE) loss [li2024rediscovering], soft intersection over union (SoftIoU) loss [rahman2016optimizing] and Dice loss [li2019dice]. To comprehensively evaluate the effectiveness of the proposed EEDM loss, we further introduce two ablation variants for comparison: the edge enhancement (EE) loss, which contains only the edge enhancement part, and the difficult mining (DM) loss, which contains only the difficult pixel mining part. This design enables a quantitative analysis of the respective contributions of edge enhancement and difficult sample mining to the overall detection performance.
From Table 2, compared with other losses, the proposed EEDM loss results in more obvious improvements in both the $IoU$ and $P_{d}$, verifying its superiority in infrared small target detection tasks. In contrast, the widely used SoftIoU loss exhibits relatively inferior performance on the complex and diverse W-IRSTD dataset. At the same time, compared with BCE loss, EE loss achieves improvements of 1.48% in the $IoU$ (from 55.44 to 56.26), 1.87% in the $P_{d}$ (from 59.35 to 60.46), and 1.67% in the Score (from 57.40 to 58.36), indicating that incorporating edge features during training can effectively enhance detection performance. In addition, compared with BCE loss, DM loss achieves improvements of 1.23% in the $IoU$ (from 55.44 to 56.12), 5.2% in the $P_{d}$ (from 59.35 to 62.45), and 3.19% in the Score (from 57.40 to 59.29), indicating the effectiveness of difficult pixel mining in this task. From Fig. 5, the ROC curve of the model trained with EEDM loss consistently stays above the others, achieving the highest AUC, which indicates superior performance across all thresholds compared with other loss functions. Furthermore, to intuitively demonstrate the performance of EEDM loss, we select edge-rich infrared small target samples and visualize the detection results using Grad-CAM [selvaraju2017grad]. Specifically, we compare the detection results obtained with the commonly used SoftIoU loss and the proposed EEDM loss. From Fig. 6, compared with the SoftIoU loss, the EEDM loss yields more precise and sensitive responses to target boundaries and fine-grained structures.
@@
Figure 6: Grad-CAM visualization comparison of detection results with SoftIoU Loss and EEDM Loss. In each group of images, the ones with green borders denote detection results using SoftIoU Loss, while the ones with red borders denote detection results using EEDM Loss. Local zoom has been applied to emphasize the differences in key areas.
2) Validation of hyperparameter stability: To comprehensively evaluate the performance of the proposed EEDM loss, we conduct detailed experiments on two key hyperparameters (the edge weight coefficient $w$ and the difficult pixel mining ratio $p$) to investigate the model’s sensitivity and robustness to different hyperparameter settings. Specifically, the edge weight coefficient $w$ is set to 1, 3, 4, 5, 7, and the difficult pixel mining ratio $p$ is set to 0.1, 0.3, 0.5, 0.7, 0.9.
From Fig. 7, the proposed EEDM loss can maintain relatively stable performance across different hyperparameter configurations. This indicates that the loss function is robust to the hyperparameters $w$ and $p$, as it consistently enhances model performance within a reasonable range of these hyperparameters. Based on the experimental results, the metric Score reaches the optimal value when $w$ is 4 and $p$ is 0.5. Specifically, when $w$ is too small, the model pays insufficient attention to edge information, leading to decreases in the $IoU$ and $P_{d}$. Conversely, when $w$ is too large, the model will overly rely on edge features, thereby neglecting semantic and other critical features, which negatively impacts the overall detection performance. It is worth noting that when $w$ is extremely reduced to 1, the EEDM loss degenerates into the DM loss. Meanwhile, when $p$ is too small, the network tends to excessively focus on difficult pixels during training, which may hinder its generalization ability and lead to degraded performance. Conversely, when $p$ is too large, the mining of difficult pixels will be insufficient, thereby weakening the model’s adaptability in complex scenarios. It is worth noting that when $p$ increases to 1, the EEDM loss degenerates into the EE loss.
Figure 7: Performance analysis of the hyperparameters $w$ and $p$ on the W-IRSTD dataset. From left to right, the figures show the model performance in terms of the $IoU$, $P_{d}$, $F_{a}$, and Score under different combinations of $(w,p)$. Red indicates the optimal value of each metric.